1. The East Street Bagel Shop has a weekly demand for 3,500 bagels, which they make at their store fresh every day. They operate 7 days a week. The raw materials and ingredients are supplied by a trusted supplier. After doing some EOQ calculations, they have decided to select the ordering lot size as 2,000 bagels worth of raw materials in every order. Additionally, after doing some risk analysis, they have also decided to keep a safety stock of 250 bagels equivalent. What is their average inventory of ingredients, in terms of equivalent bagels?
Answer #1

If the ordering lot size (Q) = 2000 bagles worth of raw materials,

Then the Average inventory = Q/2 = 2000/2 = 1000 bagles worth of raw materials